# LEMMA: Learning from Errors for MatheMatical Advancement in LLMs

### v1

@@ 1 Introduction
To address this gap, recent studies have begun exploring methods to cultivate reflection in LLMs by leveraging error data. For instance, some works Xi et al. (2024b); Li et al. (2024f); Qin et al. (2024); Guan et al. (2025) employ external critical models to critique intermediate reasoning steps or use Monte Carlo Tree Search (MCTS) to navigate complex reasoning paths and prune error branches. Others Yan et al. (2024); Han et al. (2024); Zhang et al. (2024a); Yang et al. (2024) propose self-correction frameworks that construct incorrect-correct data for fine-tuning, enabling the model to iteratively revise its outputs. However, these approaches suffer from significant limitations. MCTS-based methods introduce substantial computational overhead and complexity, while self-correction methods rely on naive and inefficient techniques to collect incorrect and correct reasoning trajectories.
@@
Experiments across mathematical reasoning benchmarks (e.g., GSM8K, MATH) demonstrate LEMMA’s effectiveness. Models fine-tuned with LEMMA achieve state-of-the-art (SOTA) performance, outperforming both standard SFT baselines and prior error-aware methods (up to 13.3% average accuracy improvement on LLaMA3-8B). LEMMA-trained models also achieve strong generalization ability through evaluation on out-of-distribution (OOD) benchmarks. Further analysis reveals LEMMA can consistently reduce the occurrence of representative error types. In contrast, while fine-tuning on the original training set (SFT) improves overall accuracy, it leads to an increase in certain error types. These results validate that structured learning from errors, guided by systematic analysis, is a powerful yet underutilized lever for advancing mathematical reasoning in LLMs.
@@ 2 Related Work
Due to the scarcity of mathematical reasoning data with detailed, human-annotated reasoning steps Song et al. (2023); Luo et al. (2023), some studies Zelikman et al. (2022); Yuan et al. (2023); Singh et al. (2024); Huang et al. (2023); Tong et al. (2024a) leverage the correct output of the model itself for fine-tuning. This strategy is also known as self-improvement or reject sampling fine-tuning.
Recently, some works Qi et al. (2024); Xi et al. (2024a); Xu et al. (2024); Xi et al. (2024b); Qin et al. (2024); Guan et al. (2025) have begun using Monte Carlo Tree Search (MCTS), Process Reward Models (PRM) Lightman et al. (2024) or critique models to further enhance self-improvement. However, these methods only utilize the correct solutions, neglecting the generated errors. Since models only learn from correct solutions, they struggle to reflect on and self-correct errors they made, leading to error accumulation Zhang et al. (2024a); Han et al. (2024); Yan et al. (2024).
@@
Data augmentation is also a prevalent strategy to enhance model performance on mathematical tasks. Magister et al. (2023) and Yue et al. (2024a) distill reasoning capabilities from LLMs into smaller LMs. Dart-Math Tong et al. (2024b) introduces a difficulty-aware answer augmentation strategy, where more solutions are generated for harder problems. To further boost model performance, several works Tang et al. (2024); Huang et al. (2024); Yue et al. (2024b); Liu et al. (2024); Wang et al. (2024a); Zhou et al. (2024); Ding et al. (2024); Lu et al. (2024a); Li et al. (2024a); Luo et al. (2023); Li et al. (2024c) synthesize more training data by creating new mathematical problems and solutions. For example, MetaMath Yu et al. (2024a) combines answer augmentation and question augmentation, as well as two backward reasoning methods Jiang et al. (2024); Weng et al. (2023), to further augment training data. Our method is orthogonal to these question augmentation methods and can be directly integrated with them.
@@
Reflection and self-correction mechanisms have been proven to be effective in enhancing the performance of large language models (LLMs) across various domains. To encourage models to identify and amend their previous errors, one common approach leverages feedback from an external verifier or critic model Shinn et al. (2024); Renze and Guven (2024); Chen et al. (2024); Li et al. (2024b, 2023); Kim et al. (2023); Wu et al. (2024); Weng et al. (2023); Du et al. (2024); Li et al. (2024d). Alternatively, some research Yang et al. (2024); Zhang et al. (2024a); Han et al. (2024); Yan et al. (2024); Qin et al. (2024); Zhang et al. (2024b); Lu et al. (2024b); Kumar et al. (2024); Singh et al. (2024) focuses on fostering the self-correction capabilities of LLMs during the generation process itself, without external feedback. To collect training data, previous works often use a relatively high temperature during the generation process to introduce errors for later critique and correction Xi et al. (2024b); Yan et al. (2024); Lu et al. (2024b); Zhang et al. (2024a); Han et al. (2024). However, research has indicated that increasing the temperature can lead to nonsensical errors or incoherent text that would not typically occur during standard generation scenarios Lu et al. (2024b); Renze (2024). Moreover, these approaches neglect different correction strategies, which could potentially restrict the model’s ability to reflect and self-correct effectively.
@@ 3.1 Task Formulation
We begin by defining key components of LEMMA. The generated reasoning trajectory $\mathbf{\tau}$ is a sequence of reasoning steps: $\mathbf{\tau}=(s_{1},s_{2},\dots,s_{n},\hat{a})$, where $\hat{a}$ is the predicted answer. A bad trajectory includes both correct steps $s^{g}$, incorrect steps $s^{g}$ and ends with an incorrect answer: $\tau^{b}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1},\dots,s^{b}_{m},\hat{a}^{b})$. Models equipped with reflection and self-correction capabilities should be able to identify and rectify the incorrect steps $\{s^{b}_{1},...,s^{b}_{m}\}$, leading to a revised trajectory. The revised trajectory can be viewed as the concatenation of a bad trajectory, a Reflection Phrase (RP), and a correct trajectory $\tau^{g}$, expressed as $\tau^{r}=\tau^{b}\oplus\textit{RP}\oplus\tau^{g}$. Here, RP represents the reflection phrases that pinpoint and correct previous errors while seamlessly transitioning to a correct step. To minimize error accumulation, the model should recognize and correct errors as early as possible. Ideally, the bad trajectory $\tau^{b}$ should be a subsequence that ends at the first erroneous step, denoted as $\tau^{b}_{sub}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1})$. The following paragraphs will detail how we collect the bad sub-trajectory $\tau^{b}_{sub}$ and the good trajectory $\tau^{g}$, ultimately constructing the error-corrective revision trajectory $\tau^{r}=\tau^{b}_{sub}\oplus\textit{RP}\oplus\tau^{g}$ for model training.
@@ 3.2 Error Analysis
To gain a holistic understanding of the mathematical reasoning errors in common LLMs, we conduct a systematic analysis on error types. We use an error taxonomy modified from Li et al. (2024e), as detailed in Tab.1. Fig.3 presents the distribution of error types for different models. Our key findings are: (1) The most common errors include “Question Misinterpretation (QM)”, “Formula Confusion Error (FC)” and “Calculation Error (CA)”. This indicates that the models require improvements in areas such as problem comprehension, formula application, and conceptual understanding. (2) The distribution of error types is relatively consistent across different models. These key insights serve as the foundation for our subsequent error-type grounded error augmentation method.
@@ 3.3 Erroneous Trajectory Collection
Based on our analysis of the relationship between error types and sampling temperature in Sec.3.2, we opt not to increase the sampling temperature. Instead, we employ a relatively low sampling temperature $t=0.7$, which is widely used in mathematical evaluation Xi et al. (2024b); Yan et al. (2024); Zhang et al. (2024b). To mitigate the reduced diversity of error steps at lower temperatures, we propose an error-type grounded mistake augmentation method that systematically generates diverse and meaningful errors for subsequent correction. Specifically, we first determine the error type distribution for each question based on our prior analysis. We then leverage a teacher model (GPT-4o) to intentionally produce erroneous trajectories given an error type, which is sampled from the previously obtained error type distribution for each question. This approach ensures that the introduced errors are both diverse and closely aligned with the error patterns observed in the student model. The prompt is provided in Fig.13 of Appendix A.4. Using this approach, we compile a comprehensive collection of bad trajectories, which consists of: (1) the erroneous trajectories generated by error augmentation and (2) those produced by the student model itself. This strategy mirrors the human learning process, where students not only reflect on and correct their own mistakes but also receive guidance from teachers, who highlight common error-prone steps based on overall performance of all students. Additionally, the teacher model annotates the first error step in each bad trajectory $\tau^{b}$. Starting from this step, the trajectory is truncated to form the bad sub-trajectory $\tau^{b}_{sub}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1})$, to minimize error accumulation.
Table 1: Error taxonomy modified from Li et al. (2024e). Corresponding examples are in Fig.12 of appendix A.2. Some infrequent error types are omitted to save space. For the full taxonomy, please check Tab.10 in the appendix.
@@ 4.1 Implementation Details
We use the training set of MATH Hendrycks et al. (2021) and GSM8K Cobbe et al. (2021) to generate the error-corrective reasoning trajectory. We utilize LLaMA3-8B to produce the self-generated errors at a temperature of $0.7$ and employ GPT-4o Hurst et al. (2024) as the teacher model to deliberately introduce errors and perform subsequent corrections. Additionally, we employ an open-source model, LLaMA-3.1-Nemotron-70B Wang et al. (2024b), as an alternative teacher model to demonstrate the generalization of our method. For each question, we generate two self-generated errors from the student model and two deliberately introduced errors from the teacher model. For LEMMA (w/ MetaMath), we collect two additional errors. One is from the student model, and the other is generated by the teacher model based on the new questions of MetaMath (Yu et al., 2024a). For each error, we apply both “Fix & Continue” and “Fresh & Restart” correction strategies once. After filtering out the trajectories with incorrect final answers, we obtain $88.90k$ error-corrective reasoning trajectories as training data. We fine-tune various base models, including general-purpose models such as LLaMA3-8B and Mistral-7B-v0.1, as well as the math-specialized model DeepSeekMath-7B and Qwen2-Math-7B. Further implementation details are available in Appendix A.3.1.
@@
For data augmentation approaches, we consider: (1) SFT: Training on the union of GSM8K and MATH training set. (2) Rejection Sampling Fine-tuning (RFT) Yuan et al. (2023): Training on the correct self-generated reasoning trajectories. (3) MetaMath Yu et al. (2024a): Combining answer augmentation, question rephrasing, and two backward reasoning methods Jiang et al. (2024); Weng et al. (2023), to augment training data. (4) GPTAug: Prompting GPT-4o to generate step-by-step solution for each question. Please refer to Appendix A.3.2 for more details regarding baseline implementation.
@@ 4.2 Main Result
(1) LEMMA significantly outperforms SOTA baseline methods across most tasks, achieving an average accuracy improvement of at least 2.3% for LLaMA3 and 4.5% for DeepSeekMath. The enhancement is particularly noticeable in challenging tasks such as MATH Hendrycks et al. (2021) and College-Math Tang et al. (2024), where LEMMA surpasses SOTA baselines by at least 4.1% and 2.8% on LLaMA3, respectively. This underscores the efficacy of reflective and self-correction capabilities for solving complex math problems. (2) Interestingly, RFT Yuan et al. (2023) lags behind all reflection and self-correction methods. We attribute this to the inherent limitation of RFT, which solely utilizes the correct self-generated solutions, forgoing the valuable opportunity to learn from failures. (3) Additionally, LEMMA demonstrates strong performance across both in-distribution and out-of-distribution datasets. While some baselines, such as MetaMath, achieve relatively good results on in-distribution datasets, they fall short compared to LEMMA on out-of-distribution datasets. Notably, scaling the data size of RefAug Zhang et al. (2024b) to $89.82k$ data (i.e., RefAug-90k) enhances in-distribution performance; however, the improvements on out-of-distribution datasets are limited or even negative.
@@ 4.3 Reflective Math Reasoning Performance
Table 3: Evaluation on reflective math reasoning using LLaMA3-8B. Best result is highlighted in bold.
@@
Following Zhang et al. (2024b), we assess the reflective reasoning abilities of LLMs fine-tuned via various methods. Tab.3 presents the results. Notably, LEMMA significantly enhances the reflective reasoning capabilities of models compared to other data augmentation methods, achieving improvements of at least 3.3% and 4.7% in accuracy on MathChat-FQA-$3rd$ and MathChat-EC, respectively. Although some data augmentation approaches, such as MetaMath, have achieved considerable performance gains on multi-turn math question answering (i.e., MathChat-FQA), they fall short in improving error correction ability, with only a 0.5% accuracy increase on MathChat-EC compared to SFT. In comparison to reflection and self-correction methods, such as ISC and RefAug, LEMMA also demonstrates notable superiority. For instance, LEMMA surpasses ISC by 4.1% and 6.2% accuracy points on MathChat-FQA-$3rd$ and MathChat-EC, respectively. Fig.10 and Fig.11 in Appendix A.2 show some output cases where the model fine-tuned with LEMMA performs reflection and self-correction to produce more accurate answers. These results further underscore LEMMA’s advantages in advancing reflective and self-correction capabilities of LLMs.

### v2

@@ 1 Introduction
To address this gap, recent studies have begun exploring methods to cultivate reflection in LLMs by leveraging error data. For instance, some works Xi et al. (2024b); Li et al. (2024e); Qin et al. (2024); Guan et al. (2025) employ external critical models to critique intermediate reasoning steps or use Monte Carlo Tree Search (MCTS) to navigate complex reasoning paths and prune error branches. Others Yan et al. (2024); Han et al. (2024); Zhang et al. (2024a); Yang et al. (2024) propose self-correction frameworks that construct incorrect-correct data for fine-tuning, enabling the model to iteratively revise its outputs. However, these approaches suffer from significant limitations. MCTS-based methods introduce substantial computational overhead and complexity, while self-correction methods rely on naive and inefficient techniques to collect incorrect and correct reasoning trajectories.
@@
Experiments across mathematical reasoning benchmarks (e.g., GSM8K, MATH) demonstrate LEMMA’s effectiveness. Models fine-tuned with LEMMA outperform both standard SFT baselines and prior error-aware methods (up to 13.3% average accuracy improvement on LLaMA3-8B). LEMMA-trained models also achieve strong generalization ability through evaluation on out-of-distribution (OOD) benchmarks. Further analysis reveals LEMMA can consistently reduce the occurrence of representative error types. In contrast, while fine-tuning on the original training set (SFT) improves overall accuracy, it leads to an increase in certain error types. These results validate that structured learning from errors, guided by systematic analysis, is a powerful yet underutilized lever for advancing mathematical reasoning in LLMs.
@@ 2 Related Work
Due to the scarcity of mathematical reasoning data with detailed, human-annotated reasoning steps Song et al. (2023); Luo et al. (2023), some studies Zelikman et al. (2022); Yuan et al. (2023); Pan et al. (2024); Singh et al. (2024); Tong et al. (2024b) leverage the correct output of the model itself for fine-tuning. This strategy is also known as self-improvement or reject sampling fine-tuning.
Recently, some works Qi et al. (2024); Xi et al. (2024a); Xu et al. (2024); Xi et al. (2024b); Qin et al. (2024); Pan et al. (2025); Guan et al. (2025) have begun using Monte Carlo Tree Search (MCTS), Process Reward Models (PRM) Lightman et al. (2024) or critique models to further enhance self-improvement. However, these methods only utilize the correct solutions, neglecting the generated errors. Since models only learn from correct solutions, they struggle to reflect on and self-correct errors they made, leading to error accumulation Zhang et al. (2024a); Han et al. (2024); Yan et al. (2024).
@@
Data augmentation is also a prevalent strategy to enhance model performance on mathematical tasks. Magister et al. (2023) and Yue et al. (2024a) distill reasoning capabilities from LLMs into smaller LMs. Dart-Math Tong et al. (2024c) introduces a difficulty-aware answer augmentation strategy, where more solutions are generated for harder problems. To further boost model performance, several works Tang et al. (2024); Huang et al. (2024); Yue et al. (2024b); Liu et al. (2024); Wang et al. (2024a); Zhou et al. (2024); Ding et al. (2024); Lu et al. (2024a); Li et al. (2024a); Luo et al. (2023); Li et al. (2024c); Lee et al. (2024) synthesize more training data by creating new mathematical problems and solutions. For example, LLM2LLM Lee et al. (2024) iteratively synthesizes more data based on the data points that the student model fails to answer. MetaMath Yu et al. (2024a) combines answer augmentation and question augmentation, as well as two backward reasoning methods Jiang et al. (2024); Weng et al. (2023), to further augment training data. Although these works have acknowledged the value of erroneous data, their focus is on sample-level augmentation through the synthesis of additional data points. In contrast, our method implements sequence-level augmentation by enriching each data point with common errors and subsequent reflections, thereby fostering a more profound capacity for self-reflection and correction. Furthermore, our method is orthogonal to these question augmentation methods and can be directly integrated with them.
@@
Reflection and self-correction mechanisms have been proven to be effective in enhancing the performance of large language models (LLMs) across various domains. To encourage models to identify and amend their previous errors, one common approach leverages feedback from an external verifier or critic model An et al. (2023); Li et al. (2023); Tong et al. (2024a); Shinn et al. (2024); Renze and Guven (2024); Li et al. (2024b); Du et al. (2024). Alternatively, some research Weng et al. (2023); Yang et al. (2024); Zhang et al. (2024a); Han et al. (2024); Yan et al. (2024); Wu et al. (2024a, b); Qin et al. (2024); Zhang et al. (2024b); Lu et al. (2024b); Kumar et al. (2024) focuses on fostering the self-correction capabilities of LLMs during the generation process itself, without external feedback. To gather training data for self-correction, prior studies introduce erroneous reasoning trajectory by employing a relatively high temperature Xi et al. (2024b); Yan et al. (2024); Zhang et al. (2024a); Han et al. (2024) or by sampling from multiple models An et al. (2023). However, recent research Lu et al. (2024b); Renze (2024) has shown that elevating the temperature can result in nonsensical errors or incoherent text that are unlikely to arise during typical generation scenarios. Additionally, sampling from multiple models An et al. (2023) can introduce irrelevant errors that the student model would not make. Moreover, these approaches neglect different correction strategies, which could potentially restrict the model’s ability to reflect and self-correction.
@@ 3.1 Task Formulation
We begin by defining key components of LEMMA. The generated reasoning trajectory $\mathbf{\tau}$ is a sequence of reasoning steps: $\mathbf{\tau}=(s_{1},s_{2},\dots,s_{n},\hat{a})$, where $\hat{a}$ is the predicted answer. A bad trajectory includes both correct steps $s^{g}$, incorrect steps $s^{g}$ and ends with an incorrect answer: $\tau^{b}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1},\dots,s^{b}_{m},\hat{a}^{b})$. Models equipped with reflection and self-correction capabilities should be able to identify and rectify the incorrect steps $\{s^{b}_{1},...,s^{b}_{m}\}$, leading to a revised trajectory. The revised trajectory can be viewed as the concatenation of a bad trajectory, a Reflection Phrase (RP), and a correct trajectory $\tau^{g}$, expressed as $\tau^{r}=\tau^{b}\oplus\textit{RP}\oplus\tau^{g}$. Here, RP represents the reflection phrases that pinpoint and correct previous errors while seamlessly transitioning to a correct step. To minimize error accumulation, the model should recognize and correct errors as early as possible. Therefore, the bad trajectory $\tau^{b}$ should be a subsequence that ends at the first erroneous step, denoted as $\tau^{b}_{sub}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1})$. The following paragraphs will detail how we collect the bad sub-trajectory $\tau^{b}_{sub}$ and the good trajectory $\tau^{g}$, ultimately constructing the error-corrective revision trajectory $\tau^{r}=\tau^{b}_{sub}\oplus\textit{RP}\oplus\tau^{g}$ for model training.
@@ 3.2 Error Analysis
To gain a holistic understanding of the mathematical reasoning errors in common LLMs, we conduct a systematic analysis on error types. We use an error taxonomy modified from Li et al. (2024d), as detailed in Tab.1. Fig.3 presents the distribution of error types for different models. Our key findings are: (1) The most common errors include “Question Misinterpretation (QM)”, “Formula Confusion Error (FC)” and “Calculation Error (CA)”. This indicates that the models require improvements in areas such as problem comprehension, formula application, and conceptual understanding. (2) The distribution of error types is relatively consistent across different models. These key insights serve as the foundation for our subsequent error-type grounded error augmentation method.
@@ 3.3 Erroneous Trajectory Collection
Based on our analysis of the relationship between error types and sampling temperature in Sec.3.2, we opt not to increase the sampling temperature. Instead, we employ a relatively low sampling temperature $t=0.7$, which is widely used in mathematical evaluation Xi et al. (2024b); Yan et al. (2024); Zhang et al. (2024b). To mitigate the reduced diversity of error steps at lower temperatures, we propose an error-type grounded mistake augmentation method that systematically generates diverse and meaningful errors for subsequent correction. Specifically, we first determine the error type distribution for each question based on our prior analysis. We then leverage a teacher model (GPT-4o) to intentionally produce erroneous trajectories given an error type, which is sampled from the previously obtained error type distribution for each question. This approach ensures that the introduced errors are both diverse and closely aligned with the errors observed using the student model.
However, as our objective is not to induce the model to learn flawed reasoning but rather to learn reflection, we also instruct the teacher model to annotate the first erroneous step within each incorrect trajectory $\tau^{b}$. Then, the incorrect trajectory is truncated at this identified step to create a sub-trajectory $\tau^{b}_{sub}=(s^{g}_{1},\dots,s^{g}_{k},s^{b}_{1})$. This design guides the fine-tuned models to timely recognize and rectify their errors, preventing them from proceeding with reasoning from an erroneous step. Using this approach, we compile a comprehensive collection of bad trajectories, which consists of: (1) the erroneous trajectories generated by error augmentation and (2) those produced by the student model itself. This strategy mirrors the human learning process, where students not only reflect on and correct their own mistakes but also receive guidance from teachers, who highlight common error-prone steps based on the overall performance of all students. This error-type grounded mistake augmentation method sets our approach apart from prior work. It avoids the inclusion of meaningless or irrelevant errors generated by using a high sampling temperature Xi et al. (2024b); Yan et al. (2024); Lu et al. (2024b); Zhang et al. (2024a); Han et al. (2024) or collected from other models An et al. (2023).
Table 1: Error taxonomy modified from Li et al. (2024d). Corresponding examples are in Fig.12 of appendix A.2. Some infrequent error types are omitted to save space. For the full taxonomy, please check Tab.12 in the appendix.
@@ 4.1 Implementation Details
We use the training set of MATH Hendrycks et al. (2021) and GSM8K Cobbe et al. (2021) to generate the error-corrective reasoning trajectory. We utilize LLaMA3-8B to produce the self-generated errors at a temperature of $0.7$ and employ GPT-4o Hurst et al. (2024) as the teacher model to deliberately introduce errors and perform subsequent corrections. Additionally, we employ an open-source model, LLaMA-3.1-Nemotron-70B Wang et al. (2024b), as an alternative teacher model to demonstrate the generalization of our method. For LEMMA (w/ MetaMath), we collect additional error-corrective reasoning trajectories based on the new questions of MetaMath (Yu et al., 2024a). For each error, we apply both “Fix & Continue” and “Fresh & Restart” correction strategies once. After filtering out the trajectories with incorrect final answers, we obtain $88.90k$ error-corrective reasoning trajectories as training data. We fine-tune various base models, including general-purpose models such as LLaMA3-8B and Mistral-7B-v0.1, as well as the math-specialized model DeepSeekMath-7B and Qwen2-Math-7B. Further implementation details are available in Appendix A.3.1.
@@
For data augmentation approaches, we consider: (1) SFT: Training on the union of GSM8K and MATH training set. (2) Rejection Sampling Fine-tuning (RFT) Yuan et al. (2023): Training on the correct self-generated reasoning trajectories. (3) MetaMath Yu et al. (2024a): Combining answer augmentation, question rephrasing, and two backward reasoning methods Jiang et al. (2024); Weng et al. (2023), to augment training data. (4) GPTAug: Prompting GPT-4o to generate step-by-step solution for each question. Appendix A.3.2 discusses more details regarding baseline implementation.
@@ 4.2 Main Result
(1) LEMMA significantly outperforms baseline methods across most tasks, achieving an average accuracy improvement of at least 2.3% for LLaMA3 and 4.5% for DeepSeekMath. The enhancement is particularly noticeable in challenging tasks such as MATH Hendrycks et al. (2021) and College-Math Tang et al. (2024), where LEMMA surpasses baselines by at least 4.1% and 2.8% on LLaMA3, respectively. This underscores the efficacy of reflective and self-correction capabilities for solving complex math problems. Note that although Dart-Math Tong et al. (2024c) achieves superior performance, our method is not directly comparable to DART-Math due to the differences in the teacher model and data size. In addition, while LEMMA lags slightly behind DART-Math on conventional math tasks, it performs better on tasks requiring reflection and follow-up reasoning, such as the follow-up QA (FQA) and Error correction (EC) tasks of MathChat Liang et al. (2024) as shown in Tab. 3. (2) Interestingly, RFT Yuan et al. (2023) lags behind all reflection and self-correction methods. We attribute this to the inherent limitation of RFT, which solely utilizes the correct self-generated solutions, forgoing the valuable opportunity to learn from failures. (3) Additionally, LEMMA demonstrates strong performance across both in-distribution and out-of-distribution datasets. While some baselines, such as MetaMath, achieve relatively good results on in-distribution datasets, they fall short compared to LEMMA on out-of-distribution datasets. Notably, scaling the data size of RefAug Zhang et al. (2024b) to $89.82k$ data (i.e., RefAug-90k) enhances in-distribution performance; however, the improvements on out-of-distribution datasets are limited or even negative.
@@ 4.3 Reflective Math Reasoning Performance
Table 3: Evaluation on reflective math reasoning using LLaMA3-8B. “DART-Math-U.” denotes DART-MATH-Uniform and “DART-Math-H.” denotes DART-MATH-Hard. Best result is highlighted in bold.
@@
Following Zhang et al. (2024b), we assess the reflective reasoning abilities of LLMs fine-tuned via various methods. Tab.3 presents the results. Notably, LEMMA significantly enhances the reflective reasoning capabilities of models compared to other data augmentation methods, achieving improvements of at least 3.3% and 4.1% in accuracy on MathChat-FQA-$3rd$ and MathChat-EC, respectively. Although some data augmentation approaches, such as MetaMath, have achieved considerable performance gains on multi-turn math question answering (i.e., MathChat-FQA), they fall short in improving error correction ability, with only a 0.5% accuracy increase on MathChat-EC compared to SFT. In comparison to reflection and self-correction methods, such as ISC and RefAug, LEMMA also demonstrates notable superiority. For instance, LEMMA surpasses ISC by 4.1% and 6.2% accuracy points on MathChat-FQA-$3rd$ and MathChat-EC, respectively. Fig.10 and Fig.11 in Appendix A.2 show some output cases where the model fine-tuned with LEMMA performs reflection and self-correction to produce more accurate answers. These results further underscore LEMMA’s advantages in advancing reflective and self-correction capabilities of LLMs.
4.4 Integration with DART-Math
Table 4: Performance comparison between different variants of LEMMA and DART-MATH on LLaMA3-8B. “DART-Math-H. (90k)” denotes a $90.7k$ dataset downsampled from DART-MATH-Hard. “LEMMA-Hard” denotes integrating LEMMA with DART-MATH.
Method Size GSM8K MATH Avg.
DART-Math-H. (90k) 90.7k 81.7 42.6 62.1
LEMMA 88.9k 79.2 38.3 58.7
DART-Math-H. (585k) 585.0k 81.1 46.6 63.8
LEMMA-Hard 90.7k 83.1 44.8 63.9
LEMMA is not contradictory but rather complementary to DART-MATH. Inspired by their work, we synthesize more error-corrective solutions for more challenging problems. Specifically, we calculate the failure rate of the LLaMA3-8B model to determine the number of solutions to synthesize for each sample. For each problem, we generate $n=f\times k_{max}$ data points, where $f$ is the failure rate and $k_{max}$ is the maximum number of solutions per question. Among these error-corrective solutions, half of the errors are generated by the student model itself, and the other half are introduced by the teacher model. We set $k_{max}$ to $20$, resulting in a total of $90.7k$ samples. We refer to this setting as LEMMA-Hard. The results presented in Tab.4 demonstrate that this setting improves the performance of our LEMMA, outperforming the model trained on a $90.7k$ dataset downsampled from DART-Math-Hard by an average accuracy gain of $1.8\%$, and is comparable to models trained on the complete $585k$ DART-Math-Hard dataset.